1.     IMC paid a dividend of $1.60 on its common stock yesterday. The dividends are expected to grow at 6% per year indefinitely. If the 10-year Treasury bond yield is 3%, IMC stock beta is 12, and market rate of return is 11%, estimate the value of IMC stock 3 years from now.


1
Expert's answer
2021-04-26T19:33:16-0400
"1.6\times(1+0.06)^3=1.9"


"12\times(1+0.11)^10=34.1"

"34.1+1.9=36"